OpenStudy (anonymous):

PLEASE HELP!!!!!! Which is an equation of a hyperbola centered at (-1, 1) with a major vertical axis of length 4 and a horizontal axis of length 3? A. (1/4)(x-1)^2 - (1/3)(y+1)^2 = 1 B. (1/16)(x-1)^2 - (1/9)(y+1)^2 = 1 C. (1/3)(x+1)^2 - (1/4)(y-1)^2 = 1 D. (1/9)(x+1)^2 - (1/16)(y-1)^2 = 1

OpenStudy (anonymous):

D. (1/9)(x+1)^2 - (1/16)(y-1)^2 = 1

OpenStudy (anonymous):

Question?

OpenStudy (anonymous):

Yes - would you be able to explain a little about what is going on, in this problem. It's kind of confusing me. Also, why is equation D. correct? Thanks! =)

OpenStudy (unklerhaukus):

lets look and option A. first,

OpenStudy (anonymous):

Do you know hyperbola standard form?

OpenStudy (anonymous):

I'm not sure. I don't remember if I've seen/used it before.

OpenStudy (anonymous):

Will you look it up first :)

OpenStudy (anonymous):

Sure

OpenStudy (unklerhaukus):

\[\frac{1}{4}(x-1)^2 - \frac13(y+1)^2 = 1\]\[\downarrow\]\[\frac{(x-1)^2}{4} - \frac{(y+1)^2}3 = 1\] \[\frac{(x-1)^2}{4} = \frac{(y+1)^2}3 + 1\]

OpenStudy (anonymous):

I took a look online, and I think this is it: x2/a2-y2/b2 = 1 There really wasn't too much info., on it.

OpenStudy (anonymous):

Actually when the center (h, k) isn't at (0,0): (x-h)^2/a2 - (y- k)^2/b2 = 1

OpenStudy (unklerhaukus):

the equation dosent make sense for some values of x for example x=1 in fact the equation dosent make sense for -1<x<3

OpenStudy (anonymous):

@UnkleRhaukus - I'm a bit confused. Is the (x-1)^2/4 = (y+1)^2/3 +1 what the formula for this particular problem would be?

OpenStudy (anonymous):

Oh, okay. So, do you think that the problem/equations are incorrectly written?

OpenStudy (unklerhaukus):

i wasent sure how to find the right answer so i was going to try to understand each potential plot and then to choose the answer when we came to it

OpenStudy (anonymous):

Oh, okay. Sure, that sounds good.

OpenStudy (unklerhaukus):

ok we should skip to D, because it look like it is the answer D. (1/9)(x+1)^2 - (1/16)(y-1)^2 = 1 \[\frac19(x+1)^2 - \frac{1}{16}(y-1)^2 = 1\] \[\frac{(x+1)^2}{9} = 1+\frac{(y-1)^2}{16} \] if x is minus 1, the LHS is zero, while the RHS is 1 + a positive number (n^2 is always positive) but \[0\neq 1+n^2\]

OpenStudy (anonymous):

Aha! I get it, now! Thank you SO much! I can't tell you how much I appreciate this!